A= 1/2h (a+b)find b
Answer by stanbon(48516)

A= (1/2)h (a+b)find b
Multiply both sides by 2 to get:
2A = h(a+b)
Divide both sides by h to get:
(2A/h) = a+b
Subtract a from both sides to get:
b = (2A/h)-a
=============
Cheers,
Stan H.